OpenStudy (anonymous):

2. What is the simplified form of (3x + 2)(4x – 3)? Use a table. 12x^2 + 18x + 6 12x^2 + x – 6 12x^2 + 18x – 6 12x^2 – x – 6

OpenStudy (anonymous):

(3x + 2)(4x – 3) 3x times 4x =12x^2 3x times -3 =-9x 2 times 4x = 8x 2 times -3 =-6 so now you have: 12x^2-9x+8x-6 then now you combine like terms -9 +8 = -x so thus you anwser is: 12x^2-x-6

OpenStudy (anonymous):

12x^2 – x – 6